OpenStudy (anonymous):

\[(x + \frac{1}{x})^2 - \frac{3}{2}(x - \frac{1}{x}) = 4\]

OpenStudy (anonymous):

just watching for fun...:)

OpenStudy (anonymous):

@moachi \[(x+\frac{1}{x})^2=(x-\frac{1}{x})^2+4\]so\[(x-\frac{1}{x})^2+4-\frac{3}{2}(x-\frac{1}{x})=4\]\[(x-\frac{1}{x})^2-\frac{3}{2}(x-\frac{1}{x})=0\]\[(x-\frac{1}{x})(x-\frac{1}{x}-\frac{3}{2})=0\]now u have 2 quadratic equation to solve\[x-\frac{1}{x}=0\]\[x-\frac{1}{x}=\frac{3}{2}\];)